# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v1

@@ Appendix A Separation of neighboring lines
The behavior of the field lines that are separated by a distance $\rho$ from an arbitrary line $\vec{x}_{0}(\ell)$ has a general form, (Boozer, 2012), as $\rho\rightarrow 0$. The trajectories of these infinitesimally separated lines are given in Courant and Snyder intrinsic coordinates by a simple Hamiltonian. This coordinate system, which is defined in Appendix B of (Courant & Snyder, 1958), is $\vec{x}(\rho,\alpha,\ell)=\rho\cos\alpha\hat{\kappa}_{0}+\rho\sin\alpha\hat{\tau}_{0}+\vec{x}_{0}(\ell)$, where the $\hat{b}_{0}\equiv\vec{B}_{0}(\ell)/B_{0}$ along the line, $d\hat{b}_{0}/d\ell=\kappa_{0}\hat{\kappa}_{0}$ is the curvature of the line, and $d\hat{\kappa}_{0}/d\ell=-(\kappa_{0}\hat{b}_{0}+\tau_{0}\hat{\tau}_{0})$ gives the torsion $\tau_{0}$. The torsion measures the extent to which the line fails to lie in a plane. Letting $\tilde{\psi}\equiv\pi B_{0}(\ell)\rho^{2}$, the neighboring trajectories are given by the Hamiltonian
@@ A.1 Maximum field-line separation
Neighboring magnetic field lines to the line $\vec{x}_{0}(\ell)$ are defined by their position at $\ell=0$, which is given by $\alpha_{0}$ and $\tilde{\psi}_{0}$. There is a maximum separation $\tilde{\psi}$, which is $\tilde{\psi}_{0}\exp(\sigma_{max}(\ell))$, and there is a minimum separation $\tilde{\psi}_{0}\exp(-\sigma_{max}(\ell))$ as a function of $\alpha_{0}$. To obtain $\sigma_{max}(\ell)$, Equation (25) can be written as $d\tilde{\psi}/d\ell=k_{e}\tilde{\psi}$, where $k_{e}\equiv 2k_{q}(\ell)\sin(2\alpha-\varphi_{q}(\ell))$ and $\alpha(\ell)$ is given by $d\alpha/d\ell=h(\alpha,\ell)$. Given a starting point, the function $k_{e}(\ell)$ can be determined with no knowledge of the dependence of $\tilde{\psi}$ on $\ell$. The exponential separation of the line is is a definite function of $\alpha_{0}$ and $\ell$ with $\sigma(\alpha_{0},\ell)=\int_{0}^{\ell}k_{e}(\ell)d\ell$. As is clear from the periodic dependence of $k_{e}$ on $\alpha$, the function $\sigma(\alpha_{0},\ell)$ will have a maximum value as a function of $\alpha_{0}$ of $\sigma_{max}(\ell)$ and a minimum value equal to minus $\sigma_{max}(\ell)$. The divergence-free nature of the magnetic field implies there must equal fluxes approaching the line $\vec{x}_{0}(\ell)$ are diverging from it.

### v2

@@ Appendix A Separation of neighboring lines
The behavior of the field lines that are separated by a distance $\rho$ from an arbitrary line $\vec{x}_{0}(\ell)$ has a general form, (Boozer:2012), as $\rho\rightarrow 0$. The trajectories of these infinitesimally separated lines are given in Courant and Snyder intrinsic coordinates by a simple Hamiltonian. This coordinate system, which is defined in Appendix B of (Courant:1958), is $\vec{x}(\rho,\alpha,\ell)=\rho\cos\alpha\hat{\kappa}_{0}+\rho\sin\alpha\hat{\tau}_{0}+\vec{x}_{0}(\ell)$, where the $\hat{b}_{0}\equiv\vec{B}_{0}(\ell)/B_{0}$ along the line, $d\hat{b}_{0}/d\ell=\kappa_{0}\hat{\kappa}_{0}$ is the curvature of the line, and $d\hat{\kappa}_{0}/d\ell=-(\kappa_{0}\hat{b}_{0}+\tau_{0}\hat{\tau}_{0})$ gives the torsion $\tau_{0}$. The torsion measures the extent to which the line fails to lie in a plane. Letting $\tilde{\psi}\equiv\pi B_{0}(\ell)\rho^{2}$, the neighboring trajectories are given by the Hamiltonian
@@ A.1 Maximum field-line separation
Neighboring magnetic field lines to the line $\vec{x}_{0}(\ell)$ are defined by their position at $\ell=0$, which is given by $\alpha_{0}$ and $\tilde{\psi}_{0}$. There is a maximum separation $\tilde{\psi}$, which is $\tilde{\psi}_{0}\exp(\sigma_{max}(\ell))$, and there is a minimum separation $\tilde{\psi}_{0}\exp(-\sigma_{max}(\ell))$ as a function of $\alpha_{0}$. To obtain $\sigma_{max}(\ell)$, Equation (26) can be written as $d\tilde{\psi}/d\ell=k_{e}\tilde{\psi}$, where $k_{e}\equiv 2k_{q}(\ell)\sin(2\alpha-\varphi_{q}(\ell))$ and $\alpha(\ell)$ is given by $d\alpha/d\ell=h(\alpha,\ell)$. Given a starting point, the function $k_{e}(\ell)$ can be determined with no knowledge of the dependence of $\tilde{\psi}$ on $\ell$. The exponential separation of the line is a definite function of $\alpha_{0}$ and $\ell$ with $\sigma(\alpha_{0},\ell)=\int_{0}^{\ell}k_{e}(\ell)d\ell$. As is clear from the periodic dependence of $k_{e}$ on $\alpha$, the function $\sigma(\alpha_{0},\ell)$ will have a maximum value as a function of $\alpha_{0}$ of $\sigma_{max}(\ell)$ and a minimum value equal to minus $\sigma_{max}(\ell)$. The divergence-free nature of the magnetic field implies there must be equal fluxes approaching the line $\vec{x}_{0}(\ell)$ as diverging from it.